OpenStudy (anonymous):

Multiply (4x – 3)^2. A. 16x^2 + 9 B. 16x^2 – 24x + 9 C. 16x^2 – 12x + 9 D. 16x^2 – 9

OpenStudy (anonymous):

Is D the correct answer?

OpenStudy (anonymous):

How?? Can you explain why D??

OpenStudy (anonymous):

\[(a-b)(a+b) = a^2 - b^2\] This is not applicable here..

OpenStudy (anonymous):

Use this here : \[(a-b)^2 = a^2 + b^2 - 2ab\]

OpenStudy (anonymous):

careful...

OpenStudy (anonymous):

\[(4x)^{2} + (-3)^{2} - 2(4x-3)\]

OpenStudy (anonymous):

is this correct?

OpenStudy (anonymous):

No..

OpenStudy (anonymous):

Last term is not right.. First two terms are looking nice to me..

OpenStudy (anonymous):

In last term you have done : \(2(a-b)\) Instead do : \(2 \cdot a \cdot b\)

OpenStudy (anonymous):

that means: a = 4x here and b = 3 here.. So \[\implies (4x)^2 + (3)^3 - 2(4x)(3)\]

OpenStudy (anonymous):

That is 3^2 there..

OpenStudy (anonymous):

\[\implies (4x)^2 + (3)^2 - 2(4x)(3)\]

OpenStudy (anonymous):

Getting RH??

OpenStudy (anonymous):

Yes yes I am sorry my connection is not working well. I understand it now. How do I distribute 2(4x)(3)

OpenStudy (anonymous):

(8x)(3) ?

OpenStudy (anonymous):

multiply the binomial (4x-3) * (4x-3)

OpenStudy (anonymous):

your answer is B

OpenStudy (anonymous):

\[16x ^{2} + 9 - (4x-3) (4x-3)\]

OpenStudy (anonymous):

like this? is this correct? @aliera95

OpenStudy (anonymous):

no.. its 16x^2-24x+9

OpenStudy (anonymous):

can you tell me how you got that answer??? I can't seem to get it:(

OpenStudy (anonymous):

since the binomial is being squared you multiply \[(4x-3) \times (4x-3) \] you then refer to the foil method. multiply the first term of the first binomial that is 4x and multiply it by the 1st and 2nd terms of the 2nd binomial.

OpenStudy (anonymous):

do the same method for the 2nd term of the 1st binomial. multiply -3 to the 1st and 2nd terms of the 2nd binomial

OpenStudy (anonymous):

finally do common like terms

OpenStudy (anonymous):

I got B as the answer